# Entering Real Social World! Benchmarking the Social Intelligence of Large Language Models from a First-person Perspective

### v2

@@ Abstract
Social intelligence is built upon three foundational pillars: cognitive intelligence, situational intelligence, and behavioral intelligence. As large language models (LLMs) become increasingly integrated into our social lives, understanding, evaluating, and developing their social intelligence are becoming increasingly important. While multiple existing works have investigated the social intelligence of LLMs, (1) most focus on a specific aspect, and the social intelligence of LLMs has yet to be systematically organized and studied; (2) position LLMs as passive observers from a third-person perspective, such as in Theory of Mind (ToM) tests. Compared to the third-person perspective, ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios. (3) a lack of comprehensive evaluation of behavioral intelligence, with specific emphasis on incorporating critical human-machine interaction scenarios. In light of this, we present EgoSocialArena, a novel framework grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence, aimed to systematically evaluate the social intelligence of LLMs from a first-person perspective. With EgoSocialArena, we conduct a comprehensive evaluation of eight prominent foundation models, even the most advanced LLMs like o1-preview lag behind human performance by 11.0 points. Our code and data are available at https://github.com/gyhou123/EgoSocialArena.
@@ 1 Introduction
In this paper, we present EgoSocialArena, a novel framework designed to systematically evaluate the social intelligence of LLMs from a first-person perspective. The development of EgoSocialArena is grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence: (1) Cognitive Intelligence: we propose a complete and generalizable workflow to transform existing static third-person ToM benchmarks into a first-person perspective. Additionally, we have newly developed a dynamic cognitive assessment in multi-turn interactive scenarios. (2) Situational Intelligence: Imagine an LLM-based Agent entering our social world - how would it respond emotionally when receiving praise or gifts [^1] This might be related to self-awareness, but the focus could be shifted more towards the application situations.? We have newly developed an assessment for such real-world social situations. Additionally, we have also developed assessments for counterfactual situations and parallel world situations. (3) Behavioral Intelligence: we incorporate existing cooperative and adversarial game environments, as well as social goal-driven interactive dialogue environments, to comprehensively evaluate the behavioral intelligence of LLMs. Overall, as illustrated in Figure 2, EgoSocialArena encompasses the evaluation of cognitive, situational, and behavioral intelligence, with eight scenarios: static cognition, dynamic cognition evolution, real-world social situation, counterfactual situation, parallel world situation, cooperative game, adversarial game, and social goal-driven human-machine interactive dialogue environment, comprising a total of 2245 data entries.
@@
We conduct extensive experiments on EgoSocialArena to evaluate 8 foundational models known for their leading performance across multiple tasks and domains. This set includes five API-based models (i.e., o1-preview, GPT-4o, GPT-4-Turbo, GPT-3.5-Turbo, and claude-3-5-sonnet-20240620) and three open-source models (LLaMa-3-8B-Chat, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-Instruct). We establish a human performance baseline by engaging qualified human annotators with a college degree or higher. Our experimental results reveal several interesting and critical insights: (1) The o1-preview model achieved the highest score of 80.6 among all models, surpassing human performance in dynamic cognition and adversarial game scenarios. Nevertheless, an 11.0 gap in overall accuracy remains when compared to the human baseline, leaving plenty of room for model improvement. Our in-depth analysis reveals that the superiority of o1-preview is mainly attributed to its powerful logical reasoning and mathematical abilities (keenly uncovering deeper patterns behind the data). (2) Comparing the performance of LLaMA3-8B-Chat and LLaMA3-70B-Chat models shows that simply scaling model size does not significantly help improve the social intelligence of LLMs. (3) Compared to the third-person perspective, LLMs show significantly improved ToM reasoning ability when operating from a first-person perspective.
@@ 4.1 Experimental Setup
We evaluate a total of eight prominent foundation LLMs, including GPT-4o [^2] https://openai.com/index/hello-gpt-4o/, o1-preview [^3] https://openai.com/index/learning-to-reason-with-llms/, GPT-4-Turbo (Achiam et al., 2023), GPT-3.5-Turbo (Achiam et al., 2023), Claude-3.5-sonnet-20240620444https://www.anthropic.com/news/claude-3-5-sonnet, LLaMa-3-8B-Chat [^5] https://ai.meta.com/blog/meta-llama-3/, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-instruct-Turbo (Dubey et al., 2024). To account for the potential influence of model parameters, we specifically compare LLaMa-3-8B-Chat with LLaMa-3-70B-Chat.
To establish a human performance baseline, we recruit 10 graduate students, all of whom have received a good basic education and possess mature cognitive abilities, to complete responses to the questions in EgoSocialArena. The average accuracy of their responses will serve as the human performance baseline. No extra tutorials or examples are provided to ensure a fair comparison. In the behavioral intelligence scenario, we similarly have these students participate in Adversarial Games and Cooperative Games, recording their average performance. For Social-Goal Driven Dialogue scenario, we use the performance of human interactions with GPT-4o as the baseline, given that GPT-4o is the best-performing LLM for this task.
@@ 4.2 Evaluation Method
Methods Cognitive Intelligence Static Cognition Dynamic Cognition-G0.8A Dynamic Cogntion Third-person First-person $\Delta$ Level 1 Level 2 Level 3 Limit Texas Open-source Models LLaMa-3-8B-Chat 50.6 66.2 +15.6 0.0 0.0 0.0 48.0 LLaMa-3-70B-Chat 58.4 63.2 +4.8 10.0 20.0 10.0 38.0 LLaMa-3.1-405B-Instruct 58.0 65.8 +7.8 80.0 20.0 20.0 56.0 API-based Models Claude-3-5-Sonnet 71.0 80.5 +9.5 50.0 10.0 40.0 66.0 GPT-3.5-Turbo 45.5 51.9 +6.4 10.0 10.0 0.0 56.0 GPT-4-Turbo 55.4 69.7 +14.3 10.0 20.0 10.0 60.0 GPT-4o 64.1 71.0 +6.9 10.0 40.0 10.0 62.0 o1-preview 71.9 77.5 +5.6 90.0 90.0 90.0 72.0 Human Human Performance 90.2 90.2 0.0 90.0 86.0 73.0 82.0
Methods Situational Intelligence Behavioral Intelligence AVG Parallel World Counterfact Real-World Adversarial Cooperative Social Goal Open-source Models LLaMa-3-8B-Chat 6.7 71.0 67.2 51.3 49.7 22.5 34.8 LLaMa-3-70B-Chat 13.3 59.0 73.2 45.0 53.3 25.5 37.3 LLaMa-3.1-405B-Instruct 36.7 66.0 77.3 52.3 65.2 34.0 52.1 API-based Models Claude-3-5-Sonnet 90.0 74.0 79.8 55.0 94.8 50.5 62.8 GPT-3.5-Turbo 13.3 37.0 72.2 46.7 50.3 33.0 34.6 GPT-4-Turbo 23.3 70.0 75.7 54.7 75.6 52.0 47.4 GPT-4o 36.7 52.0 85.8 54.0 80.8 53.0 50.5 o1-preview 86.7 90.0 84.7 56.7 96.3 52.5 80.6 Human Human Performance 93.3 91.0 90.7 56.3 100.0 64.5 91.6
@@ 4.3 Main Results
As shown in Table 2, the o1-preview model achieved the highest score of 80.6 among all models, surpassing human performance in dynamic cognition and adversarial game scenarios. Nevertheless, an 11.0 gap in overall performance remains when compared to the human baseline, leaving plenty of room for model improvement. The second-best performer is the claude-3-5-sonnet model, which demonstrate impressive results in the static cognition and parallel world scenarios. The GPT-4o model performed well in the Real Social World Situation and Social Goal-Driven interactive dialogue scenarios, likely due to being trained with a substantial amount of human feedback. Overall, the performance of open-source models lags significantly behind that of API-based models and most models still exhibit a large performance gap compared to humans. For instance, the LLaMa-3-8B-Chat model achieved an overall score of 34.8, significantly lower than the human performance of 91.6.
@@ 5 Conclusion
In this paper, considering the social intelligence of LLMs has yet to be systematically organized and studied, ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios, incorporating human-machine interaction scenario evaluation is critical and the natural approach of observing and understanding the world from an ego-centric first-person perspective for both humans and LLM-based agents, we propose the EgoToMArena framework. This framework is grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence, with eight scenarios: static, dynamic cognition; real-world, counterfactual, parallel world situation; cooperative, adversarial game, and social goal-driven human-machine interactive environment, aimed to systematically evaluate the social intelligence of LLMs from a first-person perspective. We conduct extensive experiments and observe some key insights regarding the future development of LLMs as well as the capabilities levels of the most advanced LLMs currently available.

### v3

# EgoSocialArena: Benchmarking the Social Intelligence of Large Language Models from a First-person Perspective

@@ Abstract
Social intelligence is built upon three foundational pillars: cognitive, situational, and behavioral intelligence. As large language models (LLMs) are increasingly integrated into our social lives, understanding, evaluating, and developing their social intelligence are becoming increasingly important. While multiple existing works have investigated the social intelligence of LLMs, (1) most focus on a specific aspect, and the social intelligence of LLMs has yet to be systematically organized and studied; (2) position LLMs as passive observers from a third-person perspective, such as in Theory of Mind (ToM) tests. Compared to the third-person perspective, ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios. (3) a lack of comprehensive evaluation of behavioral intelligence, with specific emphasis on incorporating critical human-machine interaction scenarios. In light of this, we present EgoSocialArena, a novel framework grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence, aimed to systematically evaluate the social intelligence of LLMs from a first-person perspective. Using EgoSocialArena, we conduct a comprehensive evaluation of eight prominent foundation models. Our findings show that even the most advanced LLMs, such as O1-preview, still fall significantly behind human performance [^1] https://github.com/gyhou123/EgoSocialArena.
@@ 1 Introduction
In this paper, we present EgoSocialArena, a novel framework designed to systematically evaluate the social intelligence of LLMs from a first-person perspective. The development of EgoSocialArena is grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence: (1) Cognitive Intelligence: we propose a complete and generalizable workflow to transform existing static third-person ToM benchmarks into a first-person perspective. Additionally, by constructing rule-based agents and reinforcement learning agents with stable capability levels and behavior strategies, we have newly developed a dynamic cognitive assessment in multi-turn interactive scenarios. (2) Situational Intelligence: Imagine an LLM-based Agent entering our social world - how would it respond emotionally when receiving praise or gifts [^2] This might be related to self-awareness, but the focus could be shifted more towards the application situations.? We have newly developed an assessment for such real-world social situations. Additionally, we have also developed assessments for counterfactual situations and parallel world situations. (3) Behavioral Intelligence: we incorporate existing cooperative and adversarial game environments, as well as social goal-driven interactive dialogue environments, to comprehensively evaluate the behavioral intelligence of LLMs. Overall, as illustrated in Figure 2, EgoSocialArena encompasses the evaluation of cognitive, situational, and behavioral intelligence, with eight scenarios: static cognition, dynamic cognition evolution, real-world social situation, counterfactual situation, parallel world situation, cooperative game, adversarial game, and social goal-driven human-machine interactive dialogue environment, comprising a total of 2245 data entries.
@@
We conduct extensive experiments on EgoSocialArena to evaluate 8 foundational models known for their leading performance across multiple tasks and domains. This set includes five API-based models (i.e., o1-preview, GPT-4o, GPT-4-Turbo, GPT-3.5-Turbo, and claude-3-5-sonnet-20240620) and three open-source models (LLaMa-3-8B-Chat, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-Instruct). We establish a human performance baseline by engaging qualified human annotators with a college degree or higher. Our experimental results reveal several interesting and critical insights: (1) The o1-preview model achieved the highest score of 80.6, surpassing human performance in dynamic cognition and adversarial game scenarios. Nevertheless, an 7.7 gap in overall accuracy remains when compared to the human baseline, leaving room for improvement. Our analysis reveals that the superiority of o1-preview is mainly attributed to its powerful logical reasoning and mathematical abilities (uncovering deeper patterns behind the data). (2) Comparing the performance of LLaMA3-8B and LLaMA3-70B models show that scaling model size does not significantly improve the social intelligence of LLMs. (3) Compared to the third-person perspective, LLMs show significantly improved ToM reasoning ability when operating from a first-person perspective.
@@ 4.1 Experimental Setup
We evaluate a total of eight prominent foundation LLMs, including GPT-4o [^3] https://openai.com/index/hello-gpt-4o/, o1-preview [^4] https://openai.com/index/learning-to-reason-with-llms/, GPT-4-Turbo (Achiam et al., 2023), GPT-3.5-Turbo (Achiam et al., 2023), Claude-3.5-sonnet-20240620555https://www.anthropic.com/news/claude-3-5-sonnet, LLaMa-3-8B-Chat [^6] https://ai.meta.com/blog/meta-llama-3/, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-instruct-Turbo (Dubey et al., 2024). To account for the potential influence of model parameters, we specifically compare LLaMa-3-8B-Chat with LLaMa-3-70B-Chat.
To establish a reliable human performance baseline, we recruit 50 graduate students, all of whom have received a good education and possess excellent cognitive abilities, to complete responses to the questions in EgoSocialArena. The average accuracy of their responses will serve as the human performance baseline. No extra tutorials or examples are provided to ensure a fair comparison. In the behavioral intelligence scenario, we similarly have these students participate in Adversarial Games and Cooperative Games, recording their average performance. For Social-Goal Driven Dialogue scenario, we use the performance of human interactions with GPT-4o as the baseline, given that GPT-4o is the best-performing LLM for this task.
@@ 4.2 Evaluation Method
Methods Cognitive Intelligence Static Cognition Dynamic Cognition-G0.8A Dynamic Cogntion Third-person First-person $\Delta$ Level 1 Level 2 Level 3 Limit Texas Open-source Models LLaMa-3-8B-Chat 50.6 66.2 +15.6 0.0 0.0 0.0 48.0 LLaMa-3-70B-Chat 58.4 63.2 +4.8 10.0 20.0 10.0 38.0 LLaMa-3.1-405B-Instruct 58.0 65.8 +7.8 80.0 20.0 20.0 56.0 API-based Models Claude-3-5-Sonnet 71.0 80.5 +9.5 50.0 10.0 40.0 66.0 GPT-3.5-Turbo 45.5 51.9 +6.4 10.0 10.0 0.0 56.0 GPT-4-Turbo 55.4 69.7 +14.3 10.0 20.0 10.0 60.0 GPT-4o 64.1 71.0 +6.9 10.0 40.0 10.0 62.0 o1-preview 71.9 77.5 +5.6 90.0 90.0 90.0 72.0 Human Human Performance 97.4 97.4 0.0 90.0 89.0 85.0 94.0
Methods Situational Intelligence Behavioral Intelligence AVG Parallel World Counterfact Real-World Adversarial Cooperative Social Goal Open-source Models LLaMa-3-8B-Chat 6.7 71.0 67.2 51.3 49.7 22.5 34.8 LLaMa-3-70B-Chat 13.3 59.0 73.2 45.0 53.3 25.5 37.3 LLaMa-3.1-405B-Instruct 36.7 66.0 77.3 52.3 65.2 34.0 52.1 API-based Models Claude-3-5-Sonnet 90.0 74.0 79.8 55.0 94.8 50.5 62.8 GPT-3.5-Turbo 13.3 37.0 72.2 46.7 50.3 33.0 34.6 GPT-4-Turbo 23.3 70.0 75.7 54.7 75.6 52.0 47.4 GPT-4o 36.7 52.0 85.8 54.0 80.8 53.0 50.5 o1-preview 86.7 90.0 84.7 56.7 96.3 52.5 80.6 Human Human Performance 96.7 97.0 96.3 56.6 100.0 69.0 88.3
@@ 4.3 Main Results
As shown in Table 2, the o1-preview model achieved the highest score of 80.6 among all models, surpassing human performance in dynamic cognition and adversarial game scenarios. Nevertheless, an 7.7 gap in overall performance remains when compared to the human baseline, leaving plenty of room for model improvement. The second-best performer is the claude-3-5-sonnet model, which demonstrate impressive results in the static cognition and parallel world scenarios. The GPT-4o model performed well in the Real Social World Situation and Social Goal-Driven interactive dialogue scenarios, likely due to being trained with a substantial amount of human feedback. Overall, the performance of open-source models lags significantly behind that of API-based models and most models still exhibit a large performance gap compared to humans. For instance, the LLaMa-3-8B-Chat model achieved an overall score of 34.8, significantly lower than the human performance of 88.3.
@@ 5 Conclusion
In this paper, we propose EgoSocialArena, a novel framework grounded in the three pillars of social intelligence: cognitive intelligence, situational intelligence, and behavioral intelligence, designed to systematically evaluate the social intelligence of LLMs from a first-person perspective. EgoSocialArena incorporates several unique design elements, including third-person to first-person perspective transformation, constructing rule-based agents and reinforcement learning agents with stable capabilities levels and behavior strategies for dynamic cognition assessment, considering non-standard and atypical social situations, evaluating the mental states of LLMs’ self after experiencing certain social events (this may be related to self-awareness), and exploring human-machine interaction. We conduct comprehensive experiments and observe some key insights regarding the future development of LLMs as well as the capabilities levels of the most advanced LLMs currently available.